Find the 31st term in the sequence whose general term is an = -3 + 4(n - 1).
Answer by richwmiller(17219)

an = -3 + 4(n - 1)
a31 = -3 + 4(31 - 1)
a31 = -3 + 4(30)
a31=120-3
a31=117